# Right Reward Right Time for Federated Learning

### v2

@@ II-A System Overview
II-B Learning Workflow
R3T is designed to perform model training, storage, and reward calculation and allocation in a fair and auditable manner, leveraging blockchain and FL technologies. It integrates a time-aware contract theory-based mechanism to address right-time and right-reward allocation where CLPs phenomenon exists. The R3T’s one-training-round process involves the following steps. 1) Prepare contracts. The cloud initiates a global model w{w}, designs a set of contract items, and distributes them to $N$ registered clients. 2) Sign contracts & Download global model. Each client signs a contract item to participate in the training process and downloads the global model w{w}. 3) Local training. Clients train w{w} using their private data $\mathcal{D}$. 4) Upload updates. Clients send their trained model updates, metadata (e.g., IDs, addresses, chosen contracts), and hashed trained model updates to their associated edge servers after training [^3] We assume that clients are truthful in reporting their updates to the cloud. If required, the cloud can verify their updates using the Trusted Execution Environments [zhang2020enabling] or reputation systems [yu2024mathcal]. Besides, given the traceability property of the underlying blockchain [yu2023ironforge] and incentive mechanism design (e.g., incentive compatibility) presented later in Section III, clients are encouraged to truthfully report to benefit from joining and contributing.. 5) Forward & Validate updates. Edges relay trained model updates to the cloud to minimize latency in the model aggregation process, and cross-validate hashed trained model updates and metadata to ensure data integrity and maintain consensus, storing records on the blockchain for transparency and dispute arbitration. 6) Receive & Aggregate model updates [^4] To avoid the straggler effect, we employ a synchronous learning update scheme, which has a provable convergence [chen2019round, kairouz2021advances].. The cloud aggregates them using an aggregation algorithm such as FedAvg [mcmahan2017communication]. 7) Settlement. The process iterates until the global model converges. Upon completion, smart contracts calculate and distribute rewards to clients to compensate for their contributions (i.e., time and effort). Edge servers also receive rewards for validating transactions and storing data.
II-C Time Frame, Client Type, Contract, and Strategy
@@
3) Cloud’s Contract: To motivate clients to join early, share resources, and conduct training, the cloud must offer transparent and fair contribution-reward bundles, compensating clients for their efforts. Recognizing the importance of CLPs, the cloud offers extra bounties to attract high-quality, resource-rich clients to participate actively. Consequently, each client’s gain is influenced by their chosen time of participation in the training process. The cloud has a contract set $\bm{\phi}=\{\phi_{k}\}_{k\in\mathcal{K}}$, which contains $K$ contract items corresponding to $K$ client types per training round. The contract item $\phi_{k}=\{e_{k},t_{k},R_{k}\}$ specifies the relationship between type-$k$ client’s effort, joining time, and reward. Specifically, $e_{k}$ is defined as the normalized local training effort, which can be determined jointly by the volume and the distribution of local data [10557146], and $t_{k}$ is the training time slot, which is either in CLPs or in non-CLPs. Also, $R_{k}$ is the corresponding reward for type-$k$ client paid by the cloud server for finishing the training round, and is given as follows:
@@ II-D Utilities
where $\lambda$ is a parameter adjusted by the cloud server to show its concerns for FL model performance (i.e., larger $\lambda$) or reward budget (i.e., smaller $\lambda$), $\bm{e}=\{e_{1},\dots,e_{K}\}$, $\bm{t}=\{t_{1},\dots,t_{K}\}$, and $\bm{r}=\{r_{1},\dots,r_{K}\}$. In each training round, multiple clients may select the same contract item $k$, and $R_{k}$ in (6) denotes the aggregate reward cost incurred by the cloud for all participating clients choosing item $k$ in that round; this aggregate notation keeps (6) compact while remaining consistent with the budget feasibility constraint, which is defined in the next section. Without the loss of generality, $g(\cdot)$ is defined as a concave function [kang2019incentive, zhou2019computation] with respect to the amount of effort and joining time. Specifically, with the same effort, each client can make different contributions and impacts on FL model performance at different training times.
II-E Blockchain smart contract-based R3T procedure
@@
Fig. 3 illustrates the functionality of blockchain and smart contracts, including four main operations. 1) First, clients and the cloud initiate their participation. During this process, R3T records information, such as IDs, and wallet addresses, for both clients and the cloud server to ensure that only registered participants can interact within the system. Additionally, the cloud publishes a contract set $\phi^{*}$, where $\phi^{*}$ is a set of optimal contract items determined by the cloud in Section III, directly onto the blockchain for transparency and accessibility to clients. 2) Upon completing the training process, clients upload their strategy profiles or selected contract items $\phi_{k}^{*}$ with their IDs, the location of the model weights, and hashed model weights to the blockchain by invoking contributionSubmit() function in the Contribution smart contract. 3) To enhance the scalability, minimize on-chain storage overhead, and provide robustness to single-point-of-failure attacks for the cloud server of R3T, dataStore() in the Contribution smart contract is employed. Global model weights are uploaded to private IPFS built by authorized edge servers while the resulting hash and corresponding training round number are recorded on-chain, ensuring data integrity and efficient and decentralized model management. 4) Finally, contract-based rewards $R_{k}$ (e.g., tokens or cryptocurrencies such as Ether in the Ethereum blockchain platform) are calculated and distributed through reward() function in the Reward smart contract.
@@
III R3T - The Proposed Incentive Mechanism for critical learning periods
III-A Problem Formulation
@@
  • Objectives: each type-$k$ client $n$ aims to maximize its utility $\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ expressed in (II-D).
Once $t_{k}$ and $r_{k}$ are given, we determine that the utility of a client in (II-D) is a strictly concave function with respect to its effort $e_{k}$. Hence, the optimal choice of effort $e_{k}$ for a type-$k$ client can be obtained by setting the first-order derivative of the client’s utility function with respect to $e_{k}$. Specifically,
@@
Substituting $e^{*}_{k}$ into (1), (II-D), (6), the utilities of a type-$k$ client and the cloud server are rewritten respectively as
@@
III-B R3T Incentivization under Complete Case
@@
III-C R3T Incentivization under Incomplete Information Case
@@ 26 $w^{t}\leftarrow\sum_{k\in\mathcal{K}}{\frac{e^{*}_{k}}{\cup_{k\in\mathcal{K}}e^{*}_{k}}}w^{t}_{k}$
According to (1), $h(t_{k})=1$ and the impact of efforts in non-CLPs is not of importance as CLPs, thereby $h(t_{k^{\prime}})=h(t_{k})=1,\forall{t_{k},t_{k^{\prime}}\notin\text{CLPs}}$ . This leads to $e_{k}>e_{k^{\prime}}$ based on (11). Moreover, via (III-C) and $h(t_{k^{\prime}})=h(t_{k})$, we can conclude that $r_{k}=r_{k^{\prime}}=r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$. Next, with the proved conditions (i.e., $h(t_{k})=h(t_{k^{\prime}})\text{ and }r_{k}=r_{k^{\prime}}$), and the expression (1), we have $R_{k}>R_{k^{\prime}}$. This completes the proof. ∎
@@ IV Experimental Results
TABLE I: Experimental Parameters
@@
IV-A Experiment Setup
Simulation parameter settings. Without loss of generality, our experimental settings refer to the setups in [zhan2020learning, zhao2021incentive]. We also define the function $g(\cdot)$ by $g(\cdot)=h(t)e$. The key parameters used in the simulations are shown in Table I.
@@ IV-D Proof of Concept of R3T
Long-term performance. For different values of the parameter of Dirichlet distribution $\alpha=\{0.4,0.7\}$ in both datasets, both R3T and CFL maintain stable test accuracies after convergence, even though R3T selects a smaller number of participating clients (cf., in line 1 of Algorithm 1). Under a highly non-IID setting (i.e., $\alpha=0.1$), R3T demonstrates more sustainable performance compared to CFL. In CFL, the practice of treating clients’ contributions and efforts equally across all training rounds, especially under a significant quantity and label distribution skew, detrimentally affects the long-term performance of the FL model, regardless of subsequent efforts. Besides, although R3T shows a slight drop in accuracy after CLPs due to the reduced number of selected clients, the overall test accuracy remains stable around the highest bound. These observations highlight the significant impact of attracting high-quality clients early, saving around 200%-300% in the total training time and 5.2%-47.6% in the total number of clients to achieve a close targeted learning performance.
V Conclusion
In this work, we studied R3T, a time-aware incentive framework for FL implemented on smart contracts. By characterizing the interactions between the client time and system capabilities, efforts, joining time, and their impact on cloud utility, we proposed a contract design that incorporates the temporal factor to derive optimal incentive strategies for the cloud. R3T transparently allocates the right reward at the right time, effectively attracting the highest-quality contributions during the early training phase, which is a critical determinant of final model performance under information symmetry and asymmetry. Simulation and proof-of-concept experimental results demonstrate the superiority of R3T over incentive benchmark mechanisms in cloud utility and economic efficiency. Furthermore, R3T reduces convergence time and the requisite number of participating clients, while achieving comparable model performance.

### v3

@@ III-A System Overview
III-B Learning Workflow
R3T is designed to perform model training, storage, and reward calculation and allocation in a fair and auditable manner, leveraging blockchain and FL technologies. It integrates a time-aware contract theory-based mechanism to address right-time and right-reward allocation where CLPs phenomenon exists. The R3T’s one-training-round process involves the following steps. 1) Prepare contracts. The cloud initiates a global model w{w}, designs a set of contract items, and distributes them to $N$ registered clients. 2) Sign contracts & Download global model. Each client signs a contract item to participate in the training process and downloads the global model w{w}. 3) Local training. Clients train w{w} using their private data $\mathcal{D}$. 4) Upload updates. Clients send their trained model updates, metadata (e.g., IDs, addresses, chosen contracts), and hashed trained model updates to their associated edge servers after training [^3] We assume that clients are truthful in reporting their updates to the cloud. If required, the cloud can verify their updates using the Trusted Execution Environments [zhang2020enabling] or reputation systems [yu2024mathcal]. Besides, given the traceability property of the underlying blockchain [yu2023ironforge] and incentive mechanism design (e.g., incentive compatibility) presented later in Section IV, clients are encouraged to truthfully report to benefit from joining and contributing.. 5) Forward & Validate updates. Edges relay trained model updates to the cloud to minimize latency in the model aggregation process, and cross-validate hashed trained model updates and metadata to ensure data integrity and maintain consensus, storing records on the blockchain for transparency and dispute arbitration. 6) Receive & Aggregate model updates [^4] To avoid the straggler effect, we employ a synchronous learning update scheme, which has a provable convergence [chen2019round, kairouz2021advances].. The cloud aggregates them using an aggregation algorithm such as FedAvg [mcmahan2017communication]. 7) Settlement. The process iterates until the global model converges. Upon completion, smart contracts calculate and distribute rewards to clients to compensate for their contributions. Edge servers also receive rewards for validating transactions and storing data.
III-C Time Frame, Client Type, Contract, and Strategy
@@
3) Cloud’s Contract: To motivate clients to join early, share resources, and conduct training, the cloud must offer transparent and fair contribution-reward bundles, compensating clients for their efforts. Recognizing the importance of CLPs, the cloud offers extra bounties to attract high-quality, resource-rich clients to participate actively. Consequently, each client’s gain is influenced by their chosen time of participation in the training process. The cloud has a contract set $\bm{\phi}=\{\phi_{k}\}_{k\in\mathcal{K}}$, which contains $K$ contract items corresponding to $K$ client types per training round. The contract item $\phi_{k}=\{e_{k},t_{k},R_{k}\}$ specifies the relationship between type-$k$ client’s effort, joining time, and reward. Specifically, $e_{k}$ is defined as the normalized local training effort, which can be determined jointly by the volume and the distribution of local data [10557146], and $t_{k}$ is the training time slot, which is either in CLPs or in non-CLPs. $R_{k}$ is the corresponding reward for type-$k$ client paid by the cloud for finishing the training round, and is given as follows:
@@ III-D Utilities
where $\lambda$ is a parameter adjusted by the cloud server to show its concerns for FL model performance (i.e., larger $\lambda$) or reward budget (i.e., smaller $\lambda$), $\bm{e}=\{e_{1},\dots,e_{K}\}$, $\bm{t}=\{t_{1},\dots,t_{K}\}$, and $\bm{r}=\{r_{1},\dots,r_{K}\}$. In each training round, multiple clients may select the same contract item $k$, and $R_{k}$ in (6) denotes the aggregate reward cost incurred by the cloud for all participating clients choosing item $k$ in that round; this aggregate notation keeps (6) compact while remaining consistent with the budget feasibility constraint, which is defined in the next section. Without the loss of generality, $g(\cdot)$ is defined as a concave function [murhekar2023incentives, kang2019incentive, zhou2019computation] with respect to the amount of effort and joining time. Specifically, with the same effort, each client can make different contributions and impacts on FL model performance at different training times.
III-E Blockchain smart contract-based R3T procedure
@@
Fig. 3 illustrates the functionality of blockchain and smart contracts, including four main operations. 1) First, clients and the cloud initiate their participation. During this process, R3T records information, such as IDs, and wallet addresses, for both clients and the cloud server to ensure that only registered participants can interact within the system. Additionally, the cloud publishes a contract set $\phi^{*}$, where $\phi^{*}$ is a set of optimal contract items determined by the cloud in Section IV, directly onto the blockchain for transparency and accessibility to clients. 2) Upon completing the training process, clients upload their strategy profiles or selected contract items $\phi_{k}^{*}$ with their IDs, the location of the model weights, and hashed model weights to the blockchain by invoking contributionSubmit() function in the Contribution smart contract. 3) To enhance the scalability, minimize on-chain storage overhead, and provide robustness to single-point-of-failure attacks for the cloud server of R3T, dataStore() in the Contribution smart contract is employed. Global model weights are uploaded to private IPFS built by authorized edge servers while the resulting hash and corresponding training round number are recorded on-chain, ensuring data integrity and efficient and decentralized model management. 4) Finally, contract-based rewards $R_{k}$ (e.g., tokens or cryptocurrencies such as Ether in the Ethereum blockchain platform) are calculated and distributed through reward() function in the Reward smart contract.
@@
IV R3T - The Proposed Incentive Mechanism for critical learning periods
IV-A Problem Formulation
@@
  • Objectives: each type-$k$ client $n$ aims to maximize its utility $\mathcal{U}_{n}(e_{k},t_{k},r_{k})$ expressed in (III-D).
Once $t_{k}$ and $r_{k}$ are given, we determine that the utility of a client in (III-D) is a strictly concave function with respect to its effort $e_{k}$. Hence, the optimal choice of effort $e_{k}$ for a type-$k$ client can be obtained by setting the first-order derivative of the client’s utility function with respect to $e_{k}$. Specifically,
@@
Substituting $e^{*}_{k}$ into (1), (III-D), (6), the utilities of a type-$k$ client and the cloud server are rewritten respectively as
@@
IV-B R3T Incentivization under Complete Case
@@
IV-C R3T Incentivization under Incomplete Information Case
@@ 26 $w^{t}\leftarrow\sum_{k\in\mathcal{K}}{\frac{e^{*}_{k}}{\cup_{k\in\mathcal{K}}e^{*}_{k}}}w^{t}_{k}$
According to (1), $h(t_{k})=1$ and the impact of efforts in non-CLPs is not of importance as CLPs, thereby $h(t_{k^{\prime}})=h(t_{k})=1,\forall{t_{k},t_{k^{\prime}}\notin\text{CLPs}}$ . This leads to $e_{k}>e_{k^{\prime}}$ based on (11). Moreover, via (IV-C) and $h(t_{k^{\prime}})=h(t_{k})$, we can conclude that $r_{k}=r_{k^{\prime}}=r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$. Next, with the proved conditions (i.e., $h(t_{k})=h(t_{k^{\prime}})\text{ and }r_{k}=r_{k^{\prime}}$), and the expression (1), we have $R_{k}>R_{k^{\prime}}$. This completes the proof. ∎
@@ V Experimental Results
TABLE II: Experimental parameters
@@
V-A Experiment Setup
Simulation parameter settings. Without loss of generality, our experimental settings refer to the setups in [zhan2020learning, zhao2021incentive]. We also define the function $g(\cdot)$ by $g(\cdot)=h(t)e$. The key parameters used in the simulations are shown in Table II.
@@ V-D Proof of Concept of R3T
Long-term performance. For different values of the parameter of Dirichlet distribution $\alpha=\{0.4,0.7\}$ in both datasets, both R3T and CFL maintain stable test accuracies after convergence, even though R3T selects a smaller number of participating clients (cf., in line 1 of Algorithm 1). Under a highly non-IID setting (i.e., $\alpha=0.1$), R3T demonstrates more sustainable performance compared to CFL. In CFL, the practice of treating clients’ contributions and efforts equally across all training rounds, especially under a significant quantity and label distribution skew, detrimentally affects the long-term performance of the FL model, regardless of subsequent efforts. Besides, although R3T shows a slight drop in accuracy after CLPs due to the reduced number of selected clients, the overall test accuracy remains stable around the highest bound. These observations highlight the significant impact of attracting high-quality clients to join early, saving up to 2x-3x faster in the total training time and 5.2%-47.6% in the total number of clients to achieve a close targeted global model learning performance.
VI Conclusion
In this work, we studied R3T, a time-aware contract-theoretic incentive framework for FL, implemented on blockchain smart contracts. Motivated by the pivotal role of CLPs and challenges caused by information asymmetries, we modeled how clients’ private time and system capabilities, effort, and joining time jointly shape the cloud utility. Building on this characterization, we derived a temporally integrated optimal contract that allocates incentives to provide the right reward at the right time, thereby inducing clients to self-reveal types, join early, and contribute higher effort, particularly in CLPs, under both complete- and incomplete-information settings. Simulation and proof-of-concept experimental results demonstrate the superiority of R3T over incentive benchmark mechanisms in cloud utility and economic efficiency, while simultaneously accelerating convergence and reducing the required number of participating clients without degrading final global model performance.
In future work, we will consider the impacts of free-riding issues and adversarial behaviors on the efficiency of R3T, especially in CLPs. Besides, it would be interesting to evaluate the efficiency and scalability of R3T across different blockchain platforms integrating FL systems, where the entire system is operated without relying on the cloud server.
Acknowledgment
This research has been conducted with financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6222; the School of Computer Science and Statistics, Trinity College Dublin. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.